Tech Staff Meeting Summary 10/2/01

Budget status was discussed.

Top priorities, in order, are currently BB and SP network maintenance, DBB, ANSS loose ends, ANSS checkup visits, and PII.

NMAD node update. Agreement with G&D is ready to go but not signed. Lease with Hunter is in the mail to CERI. We will sign both when we have both. Jim will proceed with gathering estimates of what it will take to clear around the guy wires and anchors and to replace the fence. Security will be enhanced to clear enough brush and trees to provide visual to the road but this is not a priority and will likely be costly.

ANSS station status. All stations are operational and 4 need communications. Bohannon is at CUSEC today. Mitch will call Wilkinson. HDBT needs link to Autozone and a fence. Brewer is working on CBHT problem. NAIT is waiting on NWS VPN. Jim will give Mitch a CVTN IT contact to discuss options for connecting via Music room drop. Will need optical isolation.

ANSS FY02 update Total 130 strong motion stations and 12-16 NSN backbone stations are planned. 25 are planned for ANSS-MA. CERI will install 9 and provide technical advice for others. We will not have to wait until the last minute to start this year.

Fuel Man. Problem is primarily with the truck and the van. The previous procedure was Sheila gets the transfer vouchers and gives them to Jim. Jim was to collect receipts, verify charges, and say who pays for which transfer. This requires a great deal of leg work and neither Sheila nor Jim want to do it. Neither Sheila nor Jim should have to decide who pays for what how. The solution probably is to have this information on sign out. Receipts should be collected on sign in. Proper fiscal accountability does not necessarily match receipts. Clearly we need a better system. Mitch will look into it.

GPS status. GPS operational funding is now part of MAEC HD-3 combined with REIS (formerly SG-1 and SG-2). Need better way of keeping track of SOH.

I-40 status. TDOT is interested in rapid notification and has approved funding for sometime in FY02. Greg needs to re-examine the plan.

BB network status. LNXT needs a sensor. HBAR needs radio work. HICK has problems likely associated with radios. Mitch will get a printout to assess all.

SP network status. Jim has "inventory" from the summer site visit campaign. Mitch will get a printout to make sure it is current.

ET. John has a detailed list that Mitch will give to Jim. Several sites are permitted and ready for prep and installation. Trip is on hold until we have a better idea of Mac's plans. NC stations still need maintenance and hardening but are getting better. May still need the repeater at High Peak. East TN stations faired very well during John's summer site visits. CEH has been pulled. Will use parts to get MYNC back. NSN has plans for a new eastern NC but it ain't CEH.

DBB. KOSE is ready for installation. Signed Edith Agreement has been sent to Forest Service. Need to find a repeater site between NMAD and LNXT now that we can't use the top of NMAD.

Autoreset gizmos for Lantronix and Aironets are on hold until/if we decide the hanging problems are sufficiently pathological. Doesn't look like it is for the Lantronix now the k2ew is configured more intelligently.

Panda II is still the rainy day priority. The focus is cannibalizing Panda to refurbish PandaII and dump what ain't useful.

RAMP. Probably have at least two K2's that need to go back to the factory. Bunch of other stuff to do. RAMP is a low priority until we get the FTE to help.

Marion. Didn't discuss this, but it is on hold while the hosts relocate.

Gallatin. Didn't discuss this, but I'm not sure we want to do it.